OpenStudy (anonymous):

Please help with (e^x)^x*e^28=e^11x

OpenStudy (kropot72):

\[(e ^{x})^{x}\times e ^{28}=e ^{11x}\] Is this it?

OpenStudy (anonymous):

Yes

OpenStudy (kropot72):

Take logs of both sides as follows:\[x ^{2}+28=11x\] Reaarange to form a quadratic and solve: \[x ^{2}-11x+28=0\] (x - 7)(x - 4) = 0 x = 7 and x = 4

OpenStudy (anonymous):

Very smooth work there.....

OpenStudy (kropot72):

Thx citizen :)

OpenStudy (anonymous):

I got ut right then, how bout 4^-x= 1/256, and 8^x-1.=64^2x

OpenStudy (anonymous):

These are a little different. Note that the bases on both of these problems are compatible. I'll do the first one, you should be able to follow on the second.\[4^{-x}= 1/256 \implies (2^2)^{-x}=2^{-8} \implies -2x=-8 \implies x=4\]

OpenStudy (anonymous):

kropot72, not right on this one.

OpenStudy (kropot72):

Sorry. Please give medal to AnimalAin

OpenStudy (anonymous):

No sweat. Just trying to be a team player here....

OpenStudy (anonymous):

vanessam: Have you got that?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

Nice job.